An algebra teacher drove by a farmyard full of chickens and pigs. The teacher happened to notice that there were a total of 100 heads and 270 legs. How many chickens were there? How many pigs were there?
Answer by ptaylor(2198)

Let x=number of chicks
And let y=number of pigs
We are told that:
x+y=100----------------------------eq1
We are also told that:
2x+4y=270------------------------------eq2
Multiply eq1 by 2 and then subtract it from eq2 and we get:
2y=70
y=35----------------number of pigs
Substitute this into eq1
x+35=100
x=65---------------------------number of chicks
CK
35+65=100
100=100
and
4*35+2*65=270
140+130=270
270=270
Hope this helps---ptaylor